OpenStudy (anonymous):

4x^2 + 3x = -6 Solve each equation by using the Quadratic Formula

OpenStudy (gorv):

add6 on both side

OpenStudy (anonymous):

which side would you add the 6 too

OpenStudy (anonymous):

First simplify it: \(4x^2 + 3x = -6\) Add 6 to both sides: \(4x^2 + 3x + 6 = 0\) Now it's in the form of \(ax^2 + bx + c\) We can plug it into the Quadratic formula and solve: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) Here: \(a = 4\) \(b = 3\) \(c = 6\) \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) \(x = \dfrac{-3 \pm \sqrt{3^2 - 4(4)(6)}}{2(4)}\) Multiply 2 * 4: \(x = \dfrac{-3 \pm \sqrt{3^2 - 4(4)(6)}}{8}\) Simplify exponent: \(x = \dfrac{-3 \pm \sqrt{9 - 4(4)(6)}}{8}\) Multiply: \(x = \dfrac{-3 \pm \sqrt{9 - 16(6)}}{8}\) Multiply: \(x = \dfrac{-3 \pm \sqrt{9 - 96}}{8}\) Subtract 9 - 96, what do you get?

OpenStudy (anonymous):

@good_ole_southerngirl

OpenStudy (anonymous):

Can you subtract that? @good_ole_southerngirl

OpenStudy (anonymous):

-84

OpenStudy (anonymous):

** -87

OpenStudy (anonymous):

Yes, that gives us: \(x = \dfrac{-3 \pm \sqrt{-87}}{8}\) And since we can't take the square root of a negative number, there is no solutions.

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @good_ole_southerngirl

OpenStudy (anonymous):

Would you be able to help me with possibly 2 or 4 more?

OpenStudy (anonymous):

 I don't know if I did this right